OpenStudy (anonymous):

What is the value of the expression 21^0 x (5 + 7) divided by 6 - 3 ? A. - 3 B. - 1 C. 0 D. 3

OpenStudy (anonymous):

21^0 = 1 5+7= 12 6-3=3 1*12= 12 12/3= 4 4

OpenStudy (anonymous):

4 isnt any of the choses

OpenStudy (geekfromthefutur):

Okay its 0 because anything x 0 is zero

OpenStudy (anonymous):

ohhhh ty

OpenStudy (anonymous):

anything to the power of 0 is 1

OpenStudy (geekfromthefutur):

I did the math in my head and I tried to dived it and cannot dived by zero

OpenStudy (geekfromthefutur):

any more

OpenStudy (anonymous):

yea i hve few more

OpenStudy (geekfromthefutur):

>:)

OpenStudy (anonymous):

lol

OpenStudy (geekfromthefutur):

become my fan then I help u

OpenStudy (anonymous):

kk

OpenStudy (geekfromthefutur):

y do u keep making so many profiles

OpenStudy (anonymous):

because all of the others wont let me in!!

OpenStudy (anonymous):

geek how did you get 0 as an answer

OpenStudy (geekfromthefutur):

read the responses dude/girl

OpenStudy (geekfromthefutur):

Plus I've done this assessment before and got the right answer so yay

OpenStudy (anonymous):

Convert 10^8 from exponential to standard form. 0.00000001 10,000,000 100,000,000 1,000,000,000

OpenStudy (anonymous):

there is my other question

OpenStudy (anonymous):

100,000,000

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

geek what did you multiply by 0 in the equation?

OpenStudy (anonymous):

Simplify 156^0 0 1.56 156 1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

let me get some medals

OpenStudy (anonymous):

ty can u help me with a few more ??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me close this and i will give u one :)